Evaluate sec 600° + cot 315°
Answer by tenkun(12)


sec 600 + cot 315
=sec (360+240) + cot (360-45)
=sec 240 -cot 45
=sec (270-30) - cot 45
= -cosec 30 - cot 45
= -2 -1
= -3